OpenStudy (anonymous):

Please help me understand! I will medal and fan!! Solve 4(x + 1) = 4x + 4. No solution 0 All real numbers 1

jimthompson5910 (jim_thompson5910):

what does 4(x + 1) turn into when you distribute?

jimthompson5910 (jim_thompson5910):

hint: 7*(3x + 5) = 7*3x + 7*5 = 21x + 35

OpenStudy (anonymous):

4x+4x1

jimthompson5910 (jim_thompson5910):

use * to mean multiply

OpenStudy (anonymous):

4x+4*1

OpenStudy (anonymous):

yes?

jimthompson5910 (jim_thompson5910):

yes so 4(x + 1) turns into 4x+4 when you distribute

jimthompson5910 (jim_thompson5910):

meaning that 4(x + 1) = 4x + 4 turns into 4x + 4 = 4x + 4 we have the same thing on both sides

OpenStudy (anonymous):

OK so now we just solve?

jimthompson5910 (jim_thompson5910):

yes solve for x

OpenStudy (anonymous):

4x+4=4x+4 -4 -4<< Is this right?

jimthompson5910 (jim_thompson5910):

so far so good

OpenStudy (anonymous):

ok 4x+4=4x+4 -4 -4 4x=4x /4 /4 x=x, so the answer is any possible number, right?

jimthompson5910 (jim_thompson5910):

very good

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

you're welcome